OpenStudy (anonymous):

The price-demand equation for heads of lettuce is 40p-x=25 . Where is the price of a head of lettuce and is the weekly demand (in thousands) for lettuce. How can I write an expression for revenue as a function of the weekly demand for lettuce.

OpenStudy (gw2011):

Solve for x which is the demand for lettuce: 40p-x=25 40p-25=x If you know the actual price (p), you can substitute that value into x=40p-25 and you will get the amount of lettuce demanded.

OpenStudy (anonymous):

40p-x=25 p=(25+x)/40 Revenue=p*x=((25+x)/40)*x=0.615*x+0.025*x^2